Susan is having a bakery in the heart of the city and supplies special types of cheese cookies to all the retail outlets based on the demand. Her cost of making each box of cookies is $15 and she sells it at $18. The historical data shows that her demand is normally distributed with a mean of 250 boxes per day with a standard deviation of 22 boxes. All those boxes that are not picked up on the same day by the retailers are sold in her own shop at $10 the next day. What is the optimal stocking level?
Answer #1

COST = 15
SALES = 18
SALVAGE = 10

SHORTAGE = SALES - COST = 18 - 15 = 3
OVERAGE = COST - SALVAGE = 15 - 10 = 5

SERVICE LEVEL = SHORTAGE / (SHORTAGE + OVERAGE) = 3 / (3 + 5) = 0.38
Z VALUE FOR A SERVICE LEVEL OF 0.38 = -0.31

MEAN DEMAND = 250
STANDARD DEVIATION = 22

OPTIMAL STOCKING LEVEL = MEAN DEMAND + (Z * STDEV) = 250 + (-0.31 * 22) = 243

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.